OpenStudy (anonymous):

Last Question! sin^2x -1/ cos(-x)

OpenStudy (anonymous):

\[\frac{\sin^2x-1}{\cos(-x)}= \frac{-(1-\sin^2x)}{cosx}=\frac{-\cos^2x}{cosx} = -cosx\]

OpenStudy (anonymous):

Wow that's really good. Thank you!

OpenStudy (anonymous):

since \[\cos(-\theta) = \cos \theta \rightarrow \cos(-x) = cosx\]